# The geometry of Deroin–Tholozan representations

### v2

@@ 2.2. Review of Deroin–Tholozan representations
The action-angle coordinates of $\rho$ can now be obtained by measuring some angles in the triangle chain. At the fixed point $B_{i}$ of $\rho(b_{i})$, which joins two consecutive triangles in the chain, we measure an action coordinate $\beta_{i}\in(0,2\pi)$ and an angle coordinate $\gamma_{i}\in\operatorname{\mathbb{R}}/2\pi\operatorname{\mathbb{Z}}$. We get the action coordinates $\beta=(\beta_{1},\ldots,\beta_{n-3})$ by writing the first interior angle adjacent to $B_{i}$ as $\pi-\beta_{i}/2$. That means the other interior angle adjacent to $B_{i}$ is $\beta_{i}/2$. Each of the angle coordinates $\gamma=(\gamma_{1},\ldots,\gamma_{n-3})$ is the angle $\gamma_{i}$ between the two triangles that meet $B_{i}$.
@@ 3.1. The space of hyperbolic cone metrics on a smooth sphere
Consider the smooth universal coverings $\widetilde{\Sigma}_{\textnormal{sm}}\to\Sigma^{\prime}_{\textnormal{sm}}$ and $\widetilde{S}_{\textnormal{sm}}\to S^{\prime}_{\textnormal{sm}}$. For each hyperbolic cone metric $h$ on $S_{\textnormal{sm}}$, we can find a developing map: a local isometry $\operatorname{dev}_{h}\colon\widetilde{S}_{\textnormal{sm}}\to\operatorname{\mathbb{H}}$. The developing map is unique up to postcomposition by isometries of $\operatorname{\mathbb{H}}$, and it fully describes $h$. Precomposing $\operatorname{dev}_{h}$ by an isotopy of $S_{\textnormal{sm}}$, lifted to $\widetilde{S}_{\textnormal{sm}}$, has the same effect as pulling $h$ back along that isotopy. With that in mind, consider all the developing maps of all the metrics representing elements of $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})$. Postcomposition by isometries of $\operatorname{\mathbb{H}}$ and precomposition by isotopies of $S_{\textnormal{sm}}$ lifted to $\widetilde{S}_{\textnormal{sm}}$ both act on this set, and the quotient can be identified with $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})$. Following [Ngu21], we give $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})$ the quotient topology that comes from the $\mathcal{C}^{\infty}$ topology555 The compact-open $\mathcal{C}^{\infty}$ topology, or in short $C^{\infty}$ topology, on the set $\mathcal{C}^{\infty}(X,Y)$ of smooth maps between two smooth manifolds $X$ and $Y$ is generated by the following subbase. Let $f\in\mathcal{C}^{\infty}(X,Y)$, and let $(\varphi,U)$, $(\psi,V)$ be two charts on $X$ and $Y$. Let $K\subset U$ be a compact subset with $f(K)\subset V$, and let $\varepsilon>0$. Define a sub-basic neighbourhood of $f$ as the set of all $g\in\mathcal{C}^{\infty}(X,Y)$ such that $g(K)\subset V$ and such that the derivatives of $f$ and $g$ of every order are uniformly $\varepsilon$-close on $K$ [Hir94, Chap. 2]. The $\mathcal{C}^{\infty}$ topology on $\mathcal{C}^{\infty}(X,Y)$ shouldn’t be confused with the finer Whitney $\mathcal{C}^{\infty}$ topology. on the space of smooth maps $\widetilde{S}_{\textnormal{sm}}\to\operatorname{\mathbb{H}}$.
@@
Like before, we topologize $\overline{\operatorname{Hyp}}_{\alpha}(S_{\textnormal{sm}})$ using the $\mathcal{C}^{\infty}$ topology on the space of developing maps, making the inclusion $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})\hookrightarrow\overline{\operatorname{Hyp}}_{\alpha}(S_{\textnormal{sm}})$ continuous. Using Troyanov’s theorem, we can identify each fiber of the map $\tau$ with $\operatorname{Sym}_{n-3}(\operatorname{\mathbb{CP}}^{1})$—the configuration space of $n-3$ points on $\operatorname{\mathbb{CP}}^{1}$.666 This identification is not canonical, because it involves choosing a Riemann surface isomorphism $\Sigma_{\textnormal{sm}}\to\operatorname{\mathbb{CP}}^{1}$ for each complex structure on $\Sigma_{\textnormal{sm}}$. It should be possible to do the identification continuously across fibers, turning $\overline{\operatorname{Hyp}}_{\alpha}(S_{\textnormal{sm}})$ into a trivial topological bundle over $\operatorname{Teich}(\Sigma_{\textnormal{sm}})$ whose fibers are homeomorphic to $\operatorname{Sym}_{n-3}(\operatorname{\mathbb{CP}}^{1})$. However, we won’t try to prove this. It turns out that $\operatorname{Sym}_{n-3}(\operatorname{\mathbb{CP}}^{1})$ is a complex manifold biholomorphic to $\operatorname{\mathbb{CP}}^{n-3}$.
@@
If the topological claim in footnote 6 is true, then this bijection is actually a homeomorphism.
@@ 4.1. Overview
We present a way of building hyperbolic cone spheres by gluing together elementary pieces, called samosas. The gluing occurs along the lips of slits that we cut open on each samosa, reproducing a conical cut and paste procedure. The original inspiration behind this construction was Maret’s triangle chain parameterization of DT representations. We’ll see that by constructing hyperbolic cone spheres in this manner we have a good control on their holonomy. Furthermore, it is general enough in the sense that it can be used to chart the whole space of hyperbolic cone structures on a model punctured sphere.
@@ 4.2. Building blocks
A samosa is a hyperbolic sphere with three cone singularities, which we’ll call its corners. The sphere is oriented, and the corners come with a cyclic order.
By the Gauss-Bonnet theorem, the angles of the cone singularities must sum to less than $2\pi$ [^8] If the sum is equal to $2\pi$, then the samosa is flat, instead of hyperbolic. By Corollary 1 these corner angles determine the samosa uniquely, up to isometry. We can build the samosa with corner angles $\theta=(\theta_{1},\theta_{2},\theta_{3})$ by gluing two hyperbolic triangles with interior angles $\theta/2$ along their edges. This construction is the reverse of an intrinsic decomposition of a samosa. A three-punctured sphere has one simple curve, up to isotopy, between each pair of punctures [KP10, Fig. 3], so a samosa has one simple geodesic between each pair of corners [FM12, §1.2.7]. The corners and the simple geodesics between them form a topological circle, which we’ll call the samosa’s equator. The equator is oriented to match the cyclic order of the corners. Cutting along the equator decomposes the samosa into a pair of hyperbolic triangles, which are isometric—with opposite orientations—because their sides have equal lengths. We’ll call these triangles the samosa’s hemispheres. The equator runs clockwise around the northern hemisphere, and counterclockwise around the southern hemisphere. At corner $j$, each hemisphere has a vertex with interior angle $\theta_{j}/2$.
@@
A typical slit ends in the interior of a hemisphere. However, a slit can also end on the equator or at a corner. A slit starting at corner $j$ is parameterized by a complex number $(\ell e^{i\phi})^{2\pi/\theta_{j}}$, which encodes its length $\ell>0$ and its angle $\phi$ from the negative side of the equator. The angle $\phi$ is a circle-valued parameter inside $\operatorname{\mathbb{R}}/\theta_{j}\operatorname{\mathbb{Z}}$. We usually think of $\phi$ as a number in $[-\tfrac{1}{2}\theta_{j},\tfrac{1}{2}\theta_{j}]$. We call $\ell$ the slit length and $\phi$ the slit angle.
@@
$\theta_{j+1}/2$$\ell_{\text{max}}$$d_{j-1}$$j$$j+1$$j-1$
The hyperbolic law of cosines—formula (C.2) from Appendix A— expresses $\cosh(d_{j-1})$ in terms of the corner angles of the samosa. So, it’s also possible to formulate the bound of Lemma 1 in terms of the slit angle $\phi_{j}$ and the corner angles instead of $d_{j-1}$.
@@ 4.3.1. Conical cut and paste for samosas
To join two slit samosas using a conical cut and paste, we start by cutting along two slits of equal length, one on each of the two samosas. The incision runs from the origin of the slit to its end point, unzipping it into a left lip and a right lip. We then glue each left lip to the right lip on the other samosa. Doing so, we trade two slit corners of angle $\theta$ and $\theta^{\prime}$ for a cone singularity of angle $\theta+\theta^{\prime}$ at the origin of the slit and another cone singularity of angle $4\pi$ at the end of the slit. Later in Section 4.3.3, we’ll be primarily interested in the case where $\theta+\theta^{\prime}=2\pi$. In that case, the conical cut and paste procedure turns two samosas into one hyperbolic cone sphere with four fractional singularities and one whole singularity.
@@ 4.3.2. Parameterization
The combinatorics of the gluing of a collection of slit samosas is described by a genus-0 pants assembly: a bunch [^9] We have finite sets in mind, but our construction should work for a countably infinite sets too. Countability ensures that the pants assembly describes a second-countable surface. of pairs of pants with a relation that says certain pairs of cuffs should be glued together, forming a connected surface of genus 0. The glued cuffs become pants curves that form a pants decomposition of the assembled surface, and the unglued cuffs are contracted to marked points. To get a sphere with $n$ marked points, we have to glue $n-2$ pairs of pants along $n-3$ pants curves.
@@
  • (Slit condition) Once each pair of pants is realized as a slit samosa, its slit parameters specify valid slits that don’t collide, as described in and between Definitions 4 and 5, see also Lemma 1.
@@
Let $\mathcal{S}_{\alpha,\Upsilon}$ be the space of non-degenerate samosa assemblies with angle defects $\alpha$ and underlying pants assembly $\Upsilon$ [^10] As we just saw, the definition of a samosa assembly forces $\alpha$ to satisfy Condition (1). That means $\mathcal{S}_{\alpha,\Upsilon}$ is empty unless Condition (1) holds. The space $\mathcal{S}_{\alpha,\Upsilon}$ is paramaterized by the $4(n-3)$ parameters $\beta_{b},\ell_{b},\phi_{b},\phi^{\prime}_{b}$ indexed by the pants curves $b\in{\mathcal{B}}$. The range of possible values for $\beta$ is the interior of a polytope in $\operatorname{\mathbb{R}}^{n-3}$ which is determined in terms of $\alpha$ by the inequalities above and corresponds to the polytope $\Delta$ of Corollary 2. The angles $\phi_{b}$ and $\phi_{b}^{\prime}$ can take any values on their respective circles of length $\beta_{b}$ and $\beta_{b}^{\prime}$. The slit condition in the definition of samosa assemblies (Definition 7) can be formulated in terms of an interval of admissible values for $\ell_{b}$. All small enough positive values of $\ell_{b}$ give valid slits. The upper bound is a bit more tricky and depend non-trivially on $\beta,\phi,\phi^{\prime}$ (see Lemmas 1 and 4).
@@ 4.3.3. Realizing a samosa assembly
For each unglued cuff $p$, the assembled surface has a fractional singularity, with angle defect $\alpha_{p}$. The glued edges of the slits form a closed curve representing $b$—a piecewise geodesic curve of length $2\ell_{b}$ that runs through the whole singularity, with angle $2\pi$ between the in- and out-going segments, and turns by $\beta_{b}-\pi$ at the non-singular point where the corners of the samosas were glued.
@@ 4.4. Parameterizing hyperbolic cone structures
We’d like to understand the realization map as a parameterization of hyperbolic cone structures (Definition 2). For simplicity, we’ll hereafter only consider the case where $\Upsilon$ is a chained pants assembly (Definition 6). That will be enough to accomplish the original goal of understanding Maret’s action-angle coordinate system.
@@ 4.4.1. Two trigonometric lemmas
We have an isosceles triangle with two sides of length $\ell_{b}$, which meet at an interior angle of $2\pi-\beta_{b}$. The third side has length $c_{b}$, and it makes an angle of $\kappa_{b}/2$ with each of the other two sides. The hyperbolic laws of cosines—formulae (C.1) and (C.2) from Appendix A—give
@@
Plugging the value of $\cos(\beta_{b})$ into (2) leads to, after simplifications,
@@
Equation (4) and the sign of $\kappa_{b}$ determine the value of $\beta_{b}$ (see also (11)) and (5) determines the value of $\ell_{b}$, both in terms of $c_{b}$, $\kappa_{b}$.
@@
On a samosa with one or two slits, both in the same hemisphere, the slit angles are related to the lengths of the equator segments joining two corners, the slit lengths, and the length of the shortest geodesic arc joining the endpoint of a slit and an un-slit corner by the following formulae.
  • First, assume that we’re dealing with a hemisphere containing a single slit at a corner $b$. The slit angle $\phi_{b}$ is related to the length $d_{b}$ of the equator segment joining the slit corner to an un-slit corner with apex $p\in\mathcal{P}$, the slit length $\ell_{b}$, and length $\lambda_{b}$ of the shortest geodesic arc joining the endpoint of the slit to $p$ by the equation
@@
  • Now, assume that two slits lie in the same hemisphere, at two corners $a$, $b$ of respective angle $\beta_{a}$, $\beta_{b}^{\prime}$. We’ll distinguish two configurations, both shown in the figure below. In the first configuration, neither of the two geodesic arcs joining the un-slit corner to one of the slit endpoints crosses the other slit. In that case, the slit angles $\phi_{a}$, $\phi_{b}^{\prime}$ are related to $d$, $\ell$, and $\lambda$ by the equations
@@
The hyperbolic law of cosines—formula (C.2) from Appendix A—expresses the hyperbolic cosine of an equator segment in terms of the corner angles of the samosa. So, similarly as for Lemma 1, it’s possible to formulate Equations (6), (7), (8), and (9) in terms of the corner angles of the samosa rather than the length of the equator segments.
@@
All the statements are a direct application of the hyperbolic law of cosines—formula (C.1) from Appendix A. ∎
@@
Equations (6), (7), (8), and (9) determine the value of the slit angles up to a sign. The sign of each slit angle is determined by the data of the hemisphere it lies in.

### v3

@@ 2.2. Review of Deroin–Tholozan representations
The action-angle coordinates of $\rho$ can now be obtained by measuring some angles in the triangle chain. At the fixed point $B_{i}$ of $\rho(b_{i})$, which joins two consecutive triangles in the chain, we measure an action coordinate $\beta_{i}\in(0,2\pi)$ and an angle coordinate $\gamma_{i}\in\operatorname{\mathbb{R}}/2\pi\operatorname{\mathbb{Z}}$. We get the action coordinates $\beta=(\beta_{1},\ldots,\beta_{n-3})$ by writing the first interior angle adjacent to $B_{i}$ as $\pi-\beta_{i}/2$. That means the other interior angle adjacent to $B_{i}$ is $\beta_{i}/2$. Each of the angle coordinates $\gamma=(\gamma_{1},\ldots,\gamma_{n-3})$ is the angle $\gamma_{i}$ between the two triangles that meet $B_{i}$. [^4] To be precise, $\gamma_{k}$ is the angle $\angle(C_{i+1},B_{i},C_{i+2})$.
@@ 3.1. The space of hyperbolic cone metrics on a smooth sphere
Consider the smooth universal coverings $\widetilde{\Sigma}_{\textnormal{sm}}\to\Sigma^{\prime}_{\textnormal{sm}}$ and $\widetilde{S}_{\textnormal{sm}}\to S^{\prime}_{\textnormal{sm}}$. For each hyperbolic cone metric $h$ on $S_{\textnormal{sm}}$, we can find a developing map: a local isometry $\operatorname{dev}_{h}\colon\widetilde{S}_{\textnormal{sm}}\to\operatorname{\mathbb{H}}$. The developing map is unique up to postcomposition by isometries of $\operatorname{\mathbb{H}}$, and it fully describes $h$. Precomposing $\operatorname{dev}_{h}$ by an isotopy of $S_{\textnormal{sm}}$, lifted to $\widetilde{S}_{\textnormal{sm}}$, has the same effect as pulling $h$ back along that isotopy. With that in mind, consider all the developing maps of all the metrics representing elements of $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})$. Postcomposition by isometries of $\operatorname{\mathbb{H}}$ and precomposition by isotopies of $S_{\textnormal{sm}}$ lifted to $\widetilde{S}_{\textnormal{sm}}$ both act on this set, and the quotient can be identified with $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})$. Following [Ngu21], we give $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})$ the quotient topology that comes from the $\mathcal{C}^{\infty}$ topology666 The compact-open $\mathcal{C}^{\infty}$ topology, or in short $C^{\infty}$ topology, on the set $\mathcal{C}^{\infty}(X,Y)$ of smooth maps between two smooth manifolds $X$ and $Y$ is generated by the following subbase. Let $f\in\mathcal{C}^{\infty}(X,Y)$, and let $(\varphi,U)$, $(\psi,V)$ be two charts on $X$ and $Y$. Let $K\subset U$ be a compact subset with $f(K)\subset V$, and let $\varepsilon>0$. Define a sub-basic neighbourhood of $f$ as the set of all $g\in\mathcal{C}^{\infty}(X,Y)$ such that $g(K)\subset V$ and such that the derivatives of $f$ and $g$ of every order are uniformly $\varepsilon$-close on $K$ [Hir94, Chap. 2]. The $\mathcal{C}^{\infty}$ topology on $\mathcal{C}^{\infty}(X,Y)$ shouldn’t be confused with the finer Whitney $\mathcal{C}^{\infty}$ topology. on the space of smooth maps $\widetilde{S}_{\textnormal{sm}}\to\operatorname{\mathbb{H}}$.
@@
Like before, we topologize $\overline{\operatorname{Hyp}}_{\alpha}(S_{\textnormal{sm}})$ using the $\mathcal{C}^{\infty}$ topology on the space of developing maps, making the inclusion $\operatorname{Hyp}_{\alpha}(S_{\textnormal{sm}})\hookrightarrow\overline{\operatorname{Hyp}}_{\alpha}(S_{\textnormal{sm}})$ continuous. Using Troyanov’s theorem, we can identify each fiber of the map $\tau$ with $\operatorname{Sym}_{n-3}(\operatorname{\mathbb{CP}}^{1})$—the configuration space of $n-3$ points on $\operatorname{\mathbb{CP}}^{1}$.777 This identification is not canonical, because it involves choosing a Riemann surface isomorphism $\Sigma_{\textnormal{sm}}\to\operatorname{\mathbb{CP}}^{1}$ for each complex structure on $\Sigma_{\textnormal{sm}}$. It should be possible to do the identification continuously across fibers, turning $\overline{\operatorname{Hyp}}_{\alpha}(S_{\textnormal{sm}})$ into a trivial topological bundle over $\operatorname{Teich}(\Sigma_{\textnormal{sm}})$ whose fibers are homeomorphic to $\operatorname{Sym}_{n-3}(\operatorname{\mathbb{CP}}^{1})$. However, we won’t try to prove this. It turns out that $\operatorname{Sym}_{n-3}(\operatorname{\mathbb{CP}}^{1})$ is a complex manifold biholomorphic to $\operatorname{\mathbb{CP}}^{n-3}$.
@@
If the topological claim in footnote 7 is true, then this bijection is actually a homeomorphism.
@@ 4.1. Overview
We present a way of building hyperbolic cone spheres by gluing together elementary pieces, called samosas. The gluing is done along slits that we cut open on each samosa, as described in Section 4.3.1. The original inspiration behind this construction was Maret’s triangle chain parameterization of DT representations. We’ll see that by constructing hyperbolic cone spheres in this way, we get good control on their holonomy. Furthermore, it is general enough to chart the whole space of hyperbolic cone structures on a model punctured sphere.
@@ 4.2. Building blocks
A samosa is a hyperbolic sphere with three conical singularities, which we’ll call its corners. The sphere is oriented, and the corners come with a cyclic order.
By the Gauss-Bonnet theorem, the angles of the conical singularities must sum to less than $2\pi$ [^9] If the sum is equal to $2\pi$, then the samosa is flat, instead of hyperbolic. By Corollary 1, these corner angles determine the samosa uniquely, up to isometry. We can build the samosa with corner angles $\theta=(\theta_{1},\theta_{2},\theta_{3})$ by gluing two hyperbolic triangles with interior angles $\theta/2$ along their edges. This construction is the reverse of an intrinsic decomposition of a samosa. A three-punctured sphere has one simple curve, up to isotopy, between each pair of punctures [KP10, Fig. 3], so a samosa has one simple geodesic between each pair of corners [FM12, §1.2.7]. The corners and the simple geodesics between them form a topological circle, which we’ll call the samosa’s equator. The equator is oriented to match the cyclic order of the corners. Cutting along the equator decomposes the samosa into a pair of hyperbolic triangles, which are isometric—with opposite orientations—because their sides have equal lengths. We’ll call these triangles the samosa’s hemispheres. The equator runs clockwise around the northern hemisphere, and counterclockwise around the southern hemisphere. At corner $j$, each hemisphere has a vertex with interior angle $\theta_{j}/2$.
@@
A typical slit ends in the interior of a hemisphere. However, a slit can also end on the equator or at a corner. A slit starting at corner $j$ is parameterized by its length $\ell>0$ and its angle $\phi$ from the negative side of the equator. The angle $\phi$ is a circle-valued parameter inside $\operatorname{\mathbb{R}}/\theta_{j}\operatorname{\mathbb{Z}}$. We usually think of $\phi$ as a number in $[-\tfrac{1}{2}\theta_{j},\tfrac{1}{2}\theta_{j}]$. We call $\ell$ the slit length and $\phi$ the slit angle.
@@
$\phi_{j}$$\theta_{j+1}/2$$\ell_{\text{max}}$$d_{j-1}$$j$$j+1$$j-1$
The hyperbolic law of cosines—formula (C.2) from Appendix A—expresses $\cosh(d_{j-1})$ in terms of the corner angles of the samosa. So, it’s also possible to formulate the bound from Lemma 1 in terms of the slit angle $\phi_{j}$ and the corner angles instead of $d_{j-1}$.
@@ 4.3.1. Conical cut and paste for samosas
To join two slit samosas using a conical cut and paste, we start by cutting along two slits of equal length, one on each of the two samosas. The incision runs from the origin of the slit to its end point, separating it into a left lip and a right lip. We then glue each slit’s left lip to the other slit’s right lip. In doing so, we trade the conical singularities of angles $\theta$ and $\theta^{\prime}$ at the slit corners for a conical singularity of angle $\theta+\theta^{\prime}$ at the origin of the slit and another conical singularity of angle $4\pi$ at the end of the slit. Later, in Section 4.3.3, we’ll primarily be interested in the case where $\theta+\theta^{\prime}=2\pi$. In that case, the conical cut and paste procedure turns two samosas into one hyperbolic cone sphere with four fractional singularities and one whole singularity.
@@ 4.3.2. Parameterization
Combinatorially, the gluing of a collection of slit samosas is described by a genus-0 pants assembly: a bunch [^10] We have finite sets in mind, but our construction should work for a countably infinite sets too. Countability ensures that the pants assembly describes a second-countable surface. of pairs of pants with a relation that says certain pairs of cuffs should be glued together, forming a connected surface of genus 0. The glued cuffs become pants curves that form a pants decomposition of the assembled surface, and the unglued cuffs are contracted to marked points. To get a sphere with $n$ marked points, we have to glue $n-2$ pairs of pants along $n-3$ pants curves.
@@
  • (Slit condition) Once each pair of pants is realized as a slit samosa, its slit parameters specify valid slits that don’t collide, as described in and between Definitions 4 and 5. See also Lemma 1.
@@
Let $\mathcal{S}_{\alpha,\Upsilon}$ be the space of non-degenerate samosa assemblies with angle defects $\alpha$ and underlying pants assembly $\Upsilon$ [^11] As we just saw, the definition of a samosa assembly forces $\alpha$ to satisfy Condition (1). That means $\mathcal{S}_{\alpha,\Upsilon}$ is empty unless Condition (1) holds. The space $\mathcal{S}_{\alpha,\Upsilon}$ is paramaterized by the $4(n-3)$ parameters $\beta_{b},\ell_{b},\phi_{b},\phi^{\prime}_{b}$ indexed by the pants curves $b\in{\mathcal{B}}$. The range of possible values for $\beta$ is the interior of a polytope in $\operatorname{\mathbb{R}}^{n-3}$ which is determined in terms of $\alpha$ by the inequalities above and matches the polytope $\Delta$ of Corollary 2. The angles $\phi_{b}$ and $\phi_{b}^{\prime}$ can take any values on their respective circles of length $\beta_{b}$ and $\beta_{b}^{\prime}$. The slit condition in the definition of samosa assemblies (Definition 7) can be formulated in terms of an interval of admissible values for $\ell_{b}$. All small enough positive values of $\ell_{b}$ give valid slits. The maximum slit lengths depend on $\beta,\phi,\phi^{\prime}$. Lemma 1 gives the necessary and sufficient length bound for a samosa with one slit, and Lemma 4 gives a sufficient length bound for a samosa with two slits.
@@ 4.3.3. Realizing a samosa assembly
For each unglued cuff $p$, the assembled surface has a fractional singularity, with angle defect $\alpha_{p}$. The glued edges of the slits associated with a pants curve $b\in\mathcal{B}$ form a closed curve representing $b$—a piecewise geodesic curve of length $2\ell_{b}$ that runs through the whole singularity, with angle $2\pi$ between the in- and out-going segments, and turns by $\beta_{b}-\pi$ at the non-singular point where the corners of the samosas were glued.
@@ 4.4. Parameterizing hyperbolic cone structures
We’d like to understand the realization map as a parameterization of hyperbolic cone structures (Definition 2). From now on, for simplicity, we’ll only consider the case where $\Upsilon$ is a chained pants assembly (Definition 6). That will be enough to accomplish the original goal of understanding Maret’s action-angle coordinate system.
@@ 4.4.1. Two trigonometric lemmas
We have an isosceles triangle with two sides of length $\ell_{b}$, which meet at an interior angle of $2\pi-\beta_{b}$. The third side has length $c_{b}$, and it makes an angle of $\kappa_{b}/2$ with each of the other two sides. The hyperbolic laws of cosines—formulas (C.1) and (C.2) from Appendix A—give
@@
Plugging the value of $\cos(\beta_{b})$ into (2) and simplifying shows that
@@
If we know $c_{b}$, $\kappa_{b}$, we can determine $\beta_{b}$ from the sign of $\kappa_{b}$ and Equation (4), and we can determine $\ell_{b}$ from Equation (5). The dependence of $\beta_{b}$ on the sign of $\kappa_{b}$ will be made explicit later, in Equation (11).
@@
On a samosa hemisphere containing one or two slits, the following formulas relate the slit angles to the slit lengths, the lengths of equator segments, and the distances from the slit endpoints to un-slit corners and to each other.
  • First, consider a hemisphere containing a single slit at a corner $b$. The slit angle $\phi_{b}$ is related to the length $d_{b}$ of the equator segment joining the slit corner to an un-slit corner with apex $p\in\mathcal{P}$, the slit length $\ell_{b}$, and length $\lambda_{b}$ of the shortest geodesic arc joining the endpoint of the slit to $p$ by the equation
@@
  • Now, consider a hemisphere containing two slits, at two corners $a$, $b$ with cone angles $\beta_{a}$, $\beta_{b}^{\prime}$. We’ll distinguish two configurations, both shown in the figure below. In the first configuration, neither of the two geodesic arcs joining the un-slit corner to one of the slit endpoints crosses the other slit. In that case, the slit angles $\phi_{a}$, $\phi_{b}^{\prime}$ are related to $d$, $\ell$, and $\lambda$ by the equations
@@
The hyperbolic law of cosines—formula (C.2) from Appendix A—expresses the hyperbolic cosine of an equator segment in terms of the corner angles of the samosa. So, similarly as for Lemma 1, it’s possible to formulate Equations (6), (7), (8), (9) in terms of the corner angles of the samosa rather than the lengths of the equator segments.
@@
Each statement is a direct application of the hyperbolic law of cosines—formula (C.1) from Appendix A. ∎
@@
Equations (6), (7), (8), (9) determine the absolute values of the slit angles. The sign of each slit angle is determined by which hemisphere it lies in.